### v1

@@ 3 Factorization of ${\mathbb{P}}(\alpha)$
Here we start with analysis of the posets of the form ${\mathbb{P}}(\alpha)$, where $\alpha$ is an uncountable ordinal. Preliminarily we note that these posets are homogeneous (clearly, ${\mathbb{P}}({\mathbb{X}})\upharpoonright A\cong{\mathbb{P}}({\mathbb{X}})$, for each $A\in{\mathbb{P}}({\mathbb{X}})$, whenever ${\mathbb{X}}$ is a relational structure) and atomless. Although the last claim follows from Theorems 3.2 and 4.10 we sketch a simple proof. By the next lemma we can consider the case when $\alpha$ is a limit ordinal. Then each $\xi<\alpha$ is of the form $\xi=\gamma+n$, where $\gamma\in\mathop{\rm Lim}\nolimits\cup\{0\}$ and $n\in\omega$, and the sets $E$ and $O$ of even and odd elements of $\alpha$ are disjoint copies of $\alpha$ (the corresponding embeddings are given by $f(\gamma+n)=\gamma+2n$ and $g(\gamma+n)=\gamma+2n+1$).
@@
Proof. Since $\delta_{1},s_{1}>0$, by Lemma 3.1 we have ${\mathbb{P}}(\alpha)\cong{\mathbb{P}}(\gamma)$, where, defining $k:=\sum_{i=1}^{n}s_{i}$
$\gamma=\omega^{\delta_{n}}s_{n}+\dots+\omega^{\delta_{1}}s_{1}=\sum_{j=1}^{k}{\mathbb{L}}_{j},$ (11)
and by induction on $k\in{\mathbb{N}}$ we prove that ${\mathbb{P}}(\gamma)\cong\prod_{i=1}^{n}\big{(}{\mathbb{P}}(\omega^{\delta_{i}})\big{)}^{s_{i}}(\cong\prod_{j=1}^{k}{\mathbb{P}}({\mathbb{L}}_{j})$), which will follow from the equality
@@
Assuming that $F:=f[L_{1}]\cap\gamma\neq\emptyset$ the order $F$ would be isomorphic to a final part of ${\mathbb{L}}_{1}\cong\omega^{\delta_{n}}$ and, by Fact 2.11(b) we would have $F\cong\omega^{\delta_{n}}$. So we would have $\gamma^{\prime}\cong F+f[\gamma]\subset\gamma\subset\gamma^{\prime}$ and, hence $\gamma^{\prime}\cong\gamma=\omega^{\delta_{n}}(s_{n}-1)+\dots+\omega^{\delta_{1}}s_{1}$, which is impossible since each ordinal has a unique representation in the Cantor normal form. Thus $f[L_{1}]\subset L_{1}$ and $f[L_{1}]\cong L_{1}\cong\omega^{\delta_{n}}$, which implies that $C_{1}:=f[L_{1}]$ is a cofinal subset of ${\mathbb{L}}_{1}$ and $C_{1}\in{\mathbb{P}}({\mathbb{L}}_{1})$. Consequently, $f[\gamma]\subset\gamma$ and, hence, $f[\gamma]\in{\mathbb{P}}(\gamma)$; so, by the induction hypothesis $f[\gamma]=\bigcup_{j=2}^{k+1}C_{j}$, where $\langle C_{2},\dots,C_{k}\rangle\in\prod_{j=2}^{k+1}{\mathbb{P}}({\mathbb{L}}_{j})$ and $f[\gamma^{\prime}]=\bigcup_{j=1}^{k+1}C_{j}$.
Thus, ${\mathbb{P}}(\alpha)\cong\prod_{i=1}^{n}({\mathbb{P}}(\omega^{\delta_{i}}))^{s_{i}}$ and the rest follows from Fact 2.2(e). $\Box$
4 Factorization of factors ${\mathbb{P}}(\omega^{\delta})$
@@
Second, since $\zeta_{1}<\kappa$, $\zeta_{1}\in\mathop{\rm Lim}\nolimits$ would imply that $\mathop{\rm cf}\nolimits(\zeta_{1})<\kappa$ and the final part $\kappa^{\xi_{1}}\zeta_{1}$ of $\delta$, which is a $\zeta_{1}$-sum of copies of $\kappa^{\xi_{1}}$ would be of cofinality $<\kappa$; this would imply that $\mathop{\rm cf}\nolimits(\delta)<\kappa$ as well. Thus $\kappa^{\xi_{1}}\zeta_{1}=\kappa^{\xi_{1}}(\gamma+m)=\kappa^{\xi_{1}}\gamma+\kappa^{\xi_{1}}+\dots+\kappa^{\xi_{1}}$, where $\gamma\in\mathop{\rm Lim}\nolimits_{0}\;\cap\;\kappa$ and $m\in{\mathbb{N}}$, so $\delta$ has a final part isomorphic to $\kappa^{\xi_{1}}$.
@@
Second, if $\lambda:=\mathop{\rm cf}\nolimits(\theta)<\kappa$, let $\langle\alpha_{\xi}:\xi<\lambda\rangle$ be an increasing sequence cofinal in $\theta$. Assuming that $\alpha_{\xi}+\kappa\geq\theta$, for some $\xi<\lambda$, since $\mathop{\rm cf}\nolimits(\theta)<\kappa=\mathop{\rm cf}\nolimits(\alpha_{\xi}+\kappa)$ we would have $\alpha_{\xi}+\kappa>\theta$, and, hence, $\theta=\alpha_{\xi}+\beta$, where $0<\beta<\kappa$, which is false. So $\alpha_{\xi}+\kappa<\theta$, for all $\xi<\lambda$, and $\theta=\sup\{\alpha_{\xi}:\xi<\lambda\}\leq\sup\{\alpha_{\xi}+\kappa:\xi<\lambda\}\leq\theta$. Defining $\delta_{\xi}=\alpha_{\xi}+\kappa$, for $\xi<\lambda$, we have $\theta=\sup\{\delta_{\xi}:\xi<\lambda\}$ and $\mathop{\rm cf}\nolimits(\delta_{\xi})=\kappa$, for all $\xi<\lambda$. Now, if $\lambda=\omega$ we have (B) and if $\lambda>\omega$ we have (C) ((C) is possible if $\kappa>\omega_{1}$).
@@
We note that if $A\in{\mathbb{P}}({\mathbb{X}})$, then ${\mathbb{P}}({\mathbb{A}})=\{C\subset A:C\in{\mathbb{P}}({\mathbb{X}})\}$.
Proposition 4.3
Let $\delta>0$ be an ordinal and $A,B\in{\mathbb{P}}(\omega^{\delta})$. Then we have
(a) ${\mathbb{P}}(\omega^{\delta})=\langle{\mathcal{I}}_{\omega^{\delta}}^{+},\subset\rangle\equiv_{forc}\langle{\mathcal{I}}_{\omega^{\delta}}^{+},\subset_{{\mathcal{I}}_{\omega^{\delta}}}\rangle\equiv_{forc}(P(\omega^{\delta})/{\mathcal{I}}_{\omega^{\delta}})^{+}$;
(b) $A\subset_{{\mathcal{I}}_{\omega^{\delta}}}B$ iff $\forall C\in{\mathbb{P}}({\mathbb{A}})\;C\cap B\in{\mathbb{P}}(\omega^{\delta})$;
(c) $A\not\perp_{\subset}B$ iff $A\not\perp_{\subset_{{\mathcal{I}}_{\omega^{\delta}}}}B$ iff $A\cap B\in{\mathbb{P}}(\omega^{\delta})$.
Proof. By Fact 2.11(f), ${\mathcal{I}}_{\omega^{\delta}}=\{I\subset\omega^{\delta}:\omega^{\delta}\not\hookrightarrow I\}$ is an ideal in $P(\omega^{\delta})$. Clearly we have ${\mathbb{P}}(\omega^{\delta})\subset{\mathcal{I}}_{\omega^{\delta}}^{+}$. Conversely, if $A\in{\mathcal{I}}_{\omega^{\delta}}^{+}$ and $C\in{\mathbb{P}}(\omega^{\delta})$, where $C\subset A$, then $\omega^{\delta}\hookrightarrow A\hookrightarrow\omega^{\delta}$, which implies that ${\mathbb{A}}\cong\omega^{\delta}$ and, hence, $A\in{\mathbb{P}}(\omega^{\delta})$. Thus ${\mathbb{P}}(\omega^{\delta})={\mathcal{I}}_{\omega^{\delta}}^{+}$. By Fact 2.2(f) we have $\mathop{\rm sq}\nolimits\langle{\mathcal{I}}_{\omega^{\delta}}^{+},\subset\rangle=(P(\omega^{\delta})/{\mathcal{I}}_{\omega^{\delta}})^{+}$ and by Fact 2.2(a), $\langle{\mathcal{I}}_{\omega^{\delta}}^{+},\subset\rangle\equiv_{forc}(P(\omega^{\delta})/{\mathcal{I}}_{\omega^{\delta}})^{+}$.
(b) By (8), $A\subset_{{\mathcal{I}}_{\omega^{\delta}}}B$ iff for each $C\in{\mathbb{P}}(\omega^{\delta})$ satisfying $C\subset A$ the set $C\cap B$ contains a copy of $\omega^{\delta}$, which by (a) means that $C\cap B\in{\mathbb{P}}(\omega^{\delta})$.
(c) follows from (a) and Fact 2.2(g). $\Box$
@@
(a) If $\mathop{\rm cf}\nolimits(\delta)=\kappa\geq\omega$ and $\delta=\lim_{\alpha\rightarrow\kappa}\delta_{\alpha}$, then $\mathop{\rm cf}\nolimits(\omega^{\delta})=\kappa$ and $\omega^{\delta}=\sum_{\alpha<\kappa}\omega^{\delta_{\alpha}}:=\sum_{\alpha<\kappa}L_{\alpha}$, where $L_{\alpha}\cong\omega^{\delta_{\alpha}}$, for $\alpha<\kappa$, and $L_{\alpha}\cap L_{\beta}=\emptyset$, for $\alpha\neq\beta$.
@@
Let $\theta:=\sum_{\alpha<\kappa}\omega^{\delta_{\alpha}}$ and w.l.o.g. suppose that $\theta=\sum_{\alpha<\kappa}L_{\alpha}$, where $L_{\alpha}\cong\omega^{\delta_{\alpha}}$, $\alpha<\kappa$, are disjoint. Clearly we have $\theta\geq\omega^{\delta_{\alpha}}$, for all $\alpha<\kappa$; so, $\theta\geq\sup\{\omega^{\delta_{\alpha}}:\alpha<\kappa\}=\omega^{\delta}$. Suppose that $\omega^{\delta}<\theta$. Then we would have $\omega^{\delta}\in L_{\alpha_{0}}$, for some $\alpha_{0}<\kappa$; let $\xi=\min L_{\alpha_{0}}$. By Fact 2.11(b), $\omega^{\delta}>\xi$ would imply that $\omega^{\delta}\hookrightarrow L_{\alpha_{0}}\cong\omega^{\delta_{\alpha_{0}}}$ which is impossible because $\delta_{\alpha_{0}}<\delta$ and, hence, $\omega^{\delta_{\alpha_{0}}}<\omega^{\delta}$. So $\omega^{\delta}=\xi$ and, hence, $\omega^{\delta}=\sum_{\alpha<\alpha_{0}}\omega^{\delta_{\alpha}}$. Now, if $\alpha_{0}=\alpha+1$ we would have $\omega^{\delta}\hookrightarrow L_{\alpha}\cong\omega^{\delta_{\alpha}}$, which is impossible, and if $\alpha_{0}$ is a limit ordinal we would have $\mathop{\rm cf}\nolimits(\omega^{\delta})\leq\mathop{\rm cf}\nolimits(\alpha_{0})\leq|\alpha_{0}|<\kappa$, which is impossible too. So, $\omega^{\delta}=\theta=\sum_{\alpha<\kappa}\omega^{\delta_{\alpha}}$.
@@
Proof. Let $\kappa=\mathop{\rm cf}\nolimits(\omega^{\delta})$ and let $\langle\beta_{\alpha}:\alpha<\kappa\rangle$ be an increasing cofinal sequence in $\omega^{\delta}$. Then $\omega^{\delta}=\sup\{\beta_{\alpha}:\alpha<\kappa\}$ and for $\alpha<\kappa$ we have $g(\beta_{\alpha})\in\bigcup_{i\in I}L_{i}$, which implies that there is a unique $i_{\alpha}\in I$ such that $g(\beta_{\alpha})\in L_{i_{\alpha}}$. Clearly, the function $h:\kappa\rightarrow I$ defined by $h(\alpha)=i_{\alpha}$ is non-decreasing and, since $|I|<\kappa=\bigcup_{i\in I}h^{-1}[\{i\}]$, by the regularity of $\kappa$ there is $i^{*}\in I$ such that $h^{-1}[\{i^{*}\}]\in[\kappa]^{\kappa}$. Let $h^{-1}[\{i^{*}\}]=\{\alpha_{\xi}:\xi<\ \kappa\}$ be an increasing enumeration. Now, if $\alpha\in\kappa\setminus\alpha_{0}$ and $\xi\in\kappa$ where $\alpha<\alpha_{\xi}$, then $i^{*}=h(\alpha_{0})\leq h(\alpha)\leq h(\alpha_{\xi})=i^{*}$ thus for each $\alpha\in[\alpha_{0},\kappa)$ we have $h(\alpha)=i^{*}$, that is $g(\beta_{\alpha})\in L_{i^{*}}$. So, if $\beta_{\alpha_{0}}\leq\theta<\omega^{\delta}$, then, by the cofinality of the sequence $\langle\beta_{\alpha}:\alpha<\kappa\rangle$ in $\omega^{\delta}$, there is $\alpha\in\kappa$ such that $\theta\leq\beta_{\alpha}$, which implies that $g(\beta_{\alpha_{0}})\leq g(\theta)\leq g(\beta_{\alpha})$. So, since $g(\beta_{\alpha_{0}}),g(\beta_{\alpha})\in L_{i^{*}}$ and $L_{i^{*}}$ is a convex suborder of $L$, we have $g(\theta)\in L_{i^{*}}$, for all $\theta\in[\beta_{\alpha_{0}},\omega^{\delta})$. By Fact 2.11(b) we have $\omega^{\delta}\cong[\beta_{\alpha_{0}},\omega^{\delta})\cong g[[\beta_{\alpha_{0}},\omega^{\delta})]\subset L_{i^{*}}$, that is $\omega^{\delta}\hookrightarrow L_{i^{*}}$. $\Box$
@@
Proof. (a) Let $C\in{\mathbb{P}}(L)$ and suppose that $C\subset\bigcup_{\alpha<\eta}L_{\alpha}$, for some $\eta<\kappa$. By Fact 2.11(b) we have $L\cong\bigcup_{\alpha\geq\eta}L_{\alpha}$ and, hence, $L+L\hookrightarrow L$, which is, by Fact 2.10(a), impossible.
(b) Let $C\in{\mathbb{P}}(L)$, let $f:L\hookrightarrow L$, where $C=f[L]=\sum_{\alpha\in\kappa}f[L_{\alpha}]$ and let $\alpha_{0},\eta_{0}\in\kappa$. By (a) there are $\alpha_{1}<\kappa$ and $\eta_{1}\geq\eta_{0}$ such that $f[L_{\alpha_{1}}]\cap L_{\eta_{1}}\neq\emptyset$ and, hence, $C_{1}:=\bigcup_{\alpha>\alpha_{1}}f[L_{\alpha}]\subset\bigcup_{\beta\geq\eta_{0}}L_{\beta}\cap C$. Since $C_{1}$ is a final part of $C$ and $C\cong L$, by Fact 2.11(b) there is $g:L\hookrightarrow\bigcup_{\beta\geq\eta_{0}}L_{\beta}\cap C$ and, hence, $g[L_{\alpha_{0}}]\subset\bigcup_{\beta\geq\eta_{0}}L_{\beta}\cap C$. If $\beta_{0}\geq\eta_{0}$, where $g[L_{\alpha_{0}+1}]\cap L_{\beta_{0}}\neq\emptyset$, then, since $g[L_{\alpha_{0}}]<g[L_{\alpha_{0}+1}]$, we have $g[L_{\alpha_{0}}]\subset\bigcup_{\eta_{0}\leq\beta\leq\beta_{0}}L_{\beta}\cap C$. Thus $L_{\alpha_{0}}\hookrightarrow\bigcup_{\beta\in K}L_{\beta}\cap C$, where $K=[\eta_{0},\beta_{0}]\in[\kappa\setminus\eta_{0}]^{<\kappa}$, and (19) is true.
Assuming (19) by recursion we define sequences $\langle K_{\alpha}:\alpha\in\kappa\rangle$ and $\langle f_{\alpha}:\alpha\in\kappa\rangle$ such that for each $\alpha,\alpha^{\prime}\in\kappa$ we have:
@@
Let $\alpha^{\prime}<\kappa$ and let $K_{\alpha}$ and $f_{\alpha}$ be defined for $\alpha<\alpha^{\prime}$ such that (i)-(iii) is true. By (i) and the regularity of $\kappa$ we have $\eta:=\sup(\bigcup_{\alpha<\alpha^{\prime}}K_{\alpha})+1<\kappa$ and, by (19), there are $K_{\alpha^{\prime}}\in[\kappa\setminus\eta]^{<\kappa}$ and $f_{\alpha^{\prime}}:L_{\alpha^{\prime}}\hookrightarrow\bigcup_{\beta\in K_{\alpha^{\prime}}}L_{\beta}\cap C$; so, the recursion works.
Let $f=\bigcup_{\alpha\in\kappa}f_{\alpha}$. By (ii) and (iii), $\alpha<\alpha^{\prime}$ implies $f_{\alpha}[L_{\alpha}]<f_{\alpha^{\prime}}[L_{\alpha^{\prime}}]$ and, hence, $f:L\hookrightarrow C$. Thus $C^{\prime}=f[L]\in{\mathbb{P}}(L)$, $C^{\prime}\subset C$, and by Proposition 4.3(a) we have $C\in{\mathbb{P}}(L)$.

### v2

@@ 3 Factorization of ${\mathbb{P}}(\alpha)$
Here we start with analysis of the posets of the form ${\mathbb{P}}(\alpha)$, where $\alpha$ is an uncountable ordinal. Clearly, these posets are homogeneous and atomless.
@@
Proof. Since $\delta_{1},s_{1}>0$, by Lemma 3.1 we have ${\mathbb{P}}(\alpha)\cong{\mathbb{P}}(\gamma)$, where, defining $k:=\sum_{i=1}^{n}s_{i}$, we have $\gamma=\omega^{\delta_{n}}s_{n}+\dots+\omega^{\delta_{1}}s_{1}=\sum_{j=1}^{k}{\mathbb{L}}_{j}$. By induction on $k\in{\mathbb{N}}$ we prove that ${\mathbb{P}}(\gamma)\cong\prod_{i=1}^{n}\big{(}{\mathbb{P}}(\omega^{\delta_{i}})\big{)}^{s_{i}}(\cong\prod_{j=1}^{k}{\mathbb{P}}({\mathbb{L}}_{j})$), which will follow from
@@
Assuming that $F:=f[L_{1}]\cap\gamma\neq\emptyset$ the order $F$ would be isomorphic to a final part of ${\mathbb{L}}_{1}\cong\omega^{\delta_{n}}$ and, by Fact 2.8(b) we would have $F\cong\omega^{\delta_{n}}$. So we would have $\gamma^{\prime}\cong F+f[\gamma]\subset\gamma\subset\gamma^{\prime}$ and, hence $\gamma^{\prime}\cong\gamma=\omega^{\delta_{n}}(s_{n}-1)+\dots+\omega^{\delta_{1}}s_{1}$, which is impossible since each ordinal has a unique representation in the Cantor normal form. Thus $f[L_{1}]\subset L_{1}$ and $f[L_{1}]\cong L_{1}\cong\omega^{\delta_{n}}$, which implies that $C_{1}:=f[L_{1}]$ is a cofinal subset of ${\mathbb{L}}_{1}$ and $C_{1}\in{\mathbb{P}}({\mathbb{L}}_{1})$. Consequently, $f[\gamma]\subset\gamma$ and, hence, $f[\gamma]\in{\mathbb{P}}(\gamma)$; so, by the induction hypothesis $f[\gamma]=\bigcup_{j=2}^{k+1}C_{j}$, where $\langle C_{2},\dots,C_{k}\rangle\in\prod_{j=2}^{k+1}{\mathbb{P}}({\mathbb{L}}_{j})$ and $f[\gamma^{\prime}]=\bigcup_{j=1}^{k+1}C_{j}$.
Thus, ${\mathbb{P}}(\alpha)\cong\prod_{i=1}^{n}({\mathbb{P}}(\omega^{\delta_{i}}))^{s_{i}}$ and the rest follows from Fact 2.1(c). $\Box$
4 Factorization of factors ${\mathbb{P}}(\omega^{\delta})$. ZFC results
@@
Second, since $\zeta_{1}<\kappa$, $\zeta_{1}\in\mathop{\rm Lim}\nolimits$ would imply that $\mathop{\rm cf}\nolimits(\zeta_{1})<\kappa$ and the final part $\kappa^{\xi_{1}}\zeta_{1}$ of $\delta$, which is a $\zeta_{1}$-sum of copies of $\kappa^{\xi_{1}}$ would be of cofinality $<\kappa$; this would imply that $\mathop{\rm cf}\nolimits(\delta)<\kappa$. Thus $\kappa^{\xi_{1}}\zeta_{1}=\kappa^{\xi_{1}}(\gamma+m)=\kappa^{\xi_{1}}\gamma+\kappa^{\xi_{1}}+\dots+\kappa^{\xi_{1}}$, where $\gamma\in\mathop{\rm Lim}\nolimits_{0}\;\cap\;\kappa$ and $m\in{\mathbb{N}}$, so $\delta$ has a final part isomorphic to $\kappa^{\xi_{1}}$.
@@
Second, if $\lambda:=\mathop{\rm cf}\nolimits(\theta)<\kappa$, let $\langle\alpha_{\xi}:\xi<\lambda\rangle$ be an increasing sequence cofinal in $\theta$. Assuming that $\alpha_{\xi}+\kappa\geq\theta$, for some $\xi<\lambda$, since $\mathop{\rm cf}\nolimits(\theta)<\kappa=\mathop{\rm cf}\nolimits(\alpha_{\xi}+\kappa)$ we would have $\alpha_{\xi}+\kappa>\theta$, and, hence, $\theta=\alpha_{\xi}+\beta$, where $0<\beta<\kappa$, which is false. So $\alpha_{\xi}+\kappa<\theta$, for $\xi<\lambda$, and $\theta=\sup\{\alpha_{\xi}:\xi<\lambda\}\leq\sup\{\alpha_{\xi}+\kappa:\xi<\lambda\}\leq\theta$. Defining $\delta_{\xi}=\alpha_{\xi}+\kappa$, for $\xi<\lambda$, we have $\theta=\sup\{\delta_{\xi}:\xi<\lambda\}$ and $\mathop{\rm cf}\nolimits(\delta_{\xi})=\kappa$, for all $\xi<\lambda$. Now, if $\lambda=\omega$ we have (B) and if $\lambda>\omega$ we have (C).
@@
Fact 4.3
If $\delta>0$ is an ordinal, then we have
(a) ${\mathbb{P}}(\omega^{\delta})=\langle{\mathcal{I}}_{\omega^{\delta}}^{+},\subset\rangle$, $\mathop{\rm sm}\nolimits({\mathbb{P}}(\omega^{\delta}))=\langle{\mathcal{I}}_{\omega^{\delta}}^{+},\subset_{{\mathcal{I}}_{\omega^{\delta}}}\rangle$ and $\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))=(P(\omega^{\delta})/{\mathcal{I}}_{\omega^{\delta}})^{+}$;
(b) If $\delta\geq\omega$, then $|\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))|\leq 2^{|\delta|}$.
Proof. (a) By Fact 2.8(f), ${\mathcal{I}}_{\omega^{\delta}}=\{I\subset\omega^{\delta}:\omega^{\delta}\not\hookrightarrow I\}$ is an ideal in $P(\omega^{\delta})$. Clearly we have ${\mathbb{P}}(\omega^{\delta})\subset{\mathcal{I}}_{\omega^{\delta}}^{+}$. Conversely, if $A\in{\mathcal{I}}_{\omega^{\delta}}^{+}$ and $C\in{\mathbb{P}}(\omega^{\delta})$, where $C\subset A$, then $\omega^{\delta}\hookrightarrow A\hookrightarrow\omega^{\delta}$, which implies that ${\mathbb{A}}\cong\omega^{\delta}$ and, hence, $A\in{\mathbb{P}}(\omega^{\delta})$. Thus ${\mathbb{P}}(\omega^{\delta})={\mathcal{I}}_{\omega^{\delta}}^{+}$. The rest follows from Fact 2.1(d).
(b) By Fact 2.9(a), $|\omega^{\delta}|=|\delta|$; so, by (a), $|\mathop{\rm sq}\nolimits({\mathbb{P}}(\omega^{\delta}))|\leq|P(\omega^{\delta})|=2^{|\delta|}$. $\Box$
@@
(a) If $\mathop{\rm cf}\nolimits(\delta)=\kappa\geq\omega$ and $\delta=\lim_{\alpha\rightarrow\kappa}\delta_{\alpha}$, then $\mathop{\rm cf}\nolimits(\omega^{\delta})=\kappa$ and $\omega^{\delta}=\sum_{\alpha<\kappa}\omega^{\delta_{\alpha}}:=\sum_{\alpha<\kappa}L_{\alpha}$, where $L_{\alpha}\cong\omega^{\delta_{\alpha}}$, for $\alpha<\kappa$, and $L_{\alpha}\cap L_{\beta}=\emptyset$, if $\alpha\neq\beta$.
@@
Let $\theta:=\sum_{\alpha<\kappa}\omega^{\delta_{\alpha}}$ and w.l.o.g. suppose that $\theta=\sum_{\alpha<\kappa}L_{\alpha}$, where $L_{\alpha}\cong\omega^{\delta_{\alpha}}$, $\alpha<\kappa$, are disjoint. Clearly we have $\theta\geq\omega^{\delta_{\alpha}}$, for all $\alpha<\kappa$; so, $\theta\geq\sup\{\omega^{\delta_{\alpha}}:\alpha<\kappa\}=\omega^{\delta}$. Suppose that $\omega^{\delta}<\theta$. Then we would have $\omega^{\delta}\in L_{\alpha_{0}}$, for some $\alpha_{0}<\kappa$; let $\xi=\min L_{\alpha_{0}}$. By Fact 2.8(b), $\omega^{\delta}>\xi$ would imply that $\omega^{\delta}\hookrightarrow L_{\alpha_{0}}\cong\omega^{\delta_{\alpha_{0}}}$ which is impossible because $\delta_{\alpha_{0}}<\delta$ and, hence, $\omega^{\delta_{\alpha_{0}}}<\omega^{\delta}$. So $\omega^{\delta}=\xi$ and, hence, $\omega^{\delta}=\sum_{\alpha<\alpha_{0}}\omega^{\delta_{\alpha}}$. Now, if $\alpha_{0}=\alpha+1$ we would have $\omega^{\delta}\hookrightarrow L_{\alpha}\cong\omega^{\delta_{\alpha}}$, which is impossible, and if $\alpha_{0}$ is a limit ordinal we would have $\mathop{\rm cf}\nolimits(\omega^{\delta})\leq\mathop{\rm cf}\nolimits(\alpha_{0})\leq|\alpha_{0}|<\kappa$, which is impossible too. So, $\omega^{\delta}=\theta=\sum_{\alpha<\kappa}\omega^{\delta_{\alpha}}$.
@@
Proof. Let $\kappa=\mathop{\rm cf}\nolimits(\omega^{\delta})$ and let $\langle\beta_{\alpha}:\alpha<\kappa\rangle$ be an increasing cofinal sequence in $\omega^{\delta}$. Then $\omega^{\delta}=\sup\{\beta_{\alpha}:\alpha<\kappa\}$ and for $\alpha<\kappa$ we have $g(\beta_{\alpha})\in\bigcup_{i\in I}L_{i}$, which implies that there is a unique $i_{\alpha}\in I$ such that $g(\beta_{\alpha})\in L_{i_{\alpha}}$. Clearly, the function $h:\kappa\rightarrow I$ defined by $h(\alpha)=i_{\alpha}$ is non-decreasing and, since $|I|<\kappa=\bigcup_{i\in I}h^{-1}[\{i\}]$, by the regularity of $\kappa$ there is $i^{*}\in I$ such that $h^{-1}[\{i^{*}\}]\in[\kappa]^{\kappa}$. Let $h^{-1}[\{i^{*}\}]=\{\alpha_{\xi}:\xi<\ \kappa\}$ be an increasing enumeration. Now, if $\alpha\in\kappa\setminus\alpha_{0}$ and $\xi\in\kappa$ where $\alpha<\alpha_{\xi}$, then $i^{*}=h(\alpha_{0})\leq h(\alpha)\leq h(\alpha_{\xi})=i^{*}$ thus for each $\alpha\in[\alpha_{0},\kappa)$ we have $h(\alpha)=i^{*}$, that is $g(\beta_{\alpha})\in L_{i^{*}}$. So, if $\beta_{\alpha_{0}}\leq\theta<\omega^{\delta}$, then, by the cofinality of the sequence $\langle\beta_{\alpha}:\alpha<\kappa\rangle$ in $\omega^{\delta}$, there is $\alpha\in\kappa$ such that $\theta\leq\beta_{\alpha}$, which implies that $g(\beta_{\alpha_{0}})\leq g(\theta)\leq g(\beta_{\alpha})$. So, since $g(\beta_{\alpha_{0}}),g(\beta_{\alpha})\in L_{i^{*}}$ and $L_{i^{*}}$ is a convex suborder of $L$, we have $g(\theta)\in L_{i^{*}}$, for all $\theta\in[\beta_{\alpha_{0}},\omega^{\delta})$. By Fact 2.8(b) we have $\omega^{\delta}\cong[\beta_{\alpha_{0}},\omega^{\delta})\cong g[[\beta_{\alpha_{0}},\omega^{\delta})]\subset L_{i^{*}}$, that is $\omega^{\delta}\hookrightarrow L_{i^{*}}$. $\Box$
@@
Proof. (a) Let $C\in{\mathbb{P}}(L)$ and suppose that $C\subset\bigcup_{\alpha<\eta}L_{\alpha}$, for some $\eta<\kappa$. By Fact 2.8(b) we have $L\cong\bigcup_{\alpha\geq\eta}L_{\alpha}$ and, hence, $L+L\hookrightarrow L$, which is, by Fact 2.7(a), impossible.
(b) Let $C\in{\mathbb{P}}(L)$, let $f:L\hookrightarrow L$, where $C=f[L]=\sum_{\alpha\in\kappa}f[L_{\alpha}]$ and let $\alpha_{0},\eta_{0}\in\kappa$. By (a) there are $\alpha_{1}<\kappa$ and $\eta_{1}\geq\eta_{0}$ such that $f[L_{\alpha_{1}}]\cap L_{\eta_{1}}\neq\emptyset$ and, hence, $C_{1}:=\bigcup_{\alpha>\alpha_{1}}f[L_{\alpha}]\subset\bigcup_{\beta\geq\eta_{0}}L_{\beta}\cap C$. Since $C_{1}$ is a final part of $C$ and $C\cong L$, by Fact 2.8(b) there is $g:L\hookrightarrow\bigcup_{\beta\geq\eta_{0}}L_{\beta}\cap C$ and, hence, $g[L_{\alpha_{0}}]\subset\bigcup_{\beta\geq\eta_{0}}L_{\beta}\cap C$. If $\beta_{0}\geq\eta_{0}$, where $g[L_{\alpha_{0}+1}]\cap L_{\beta_{0}}\neq\emptyset$, then, since $g[L_{\alpha_{0}}]<g[L_{\alpha_{0}+1}]$, we have $g[L_{\alpha_{0}}]\subset\bigcup_{\eta_{0}\leq\beta\leq\beta_{0}}L_{\beta}\cap C$. Thus $L_{\alpha_{0}}\hookrightarrow\bigcup_{\beta\in K}L_{\beta}\cap C$, where $K=[\eta_{0},\beta_{0}]\in[\kappa\setminus\eta_{0}]^{<\kappa}$, and (10) is true.
Assuming that (10) holds, by recursion we define sequences $\langle K_{\alpha}:\alpha\in\kappa\rangle$ and $\langle f_{\alpha}:\alpha\in\kappa\rangle$ such that for each $\alpha,\alpha^{\prime}\in\kappa$ we have:
@@
Let $\alpha^{\prime}<\kappa$ and let $K_{\alpha}$ and $f_{\alpha}$ be defined for $\alpha<\alpha^{\prime}$ such that (i)-(iii) is true. By (i) and the regularity of $\kappa$ we have $\eta:=\sup(\bigcup_{\alpha<\alpha^{\prime}}K_{\alpha})+1<\kappa$ and, by (10), there are $K_{\alpha^{\prime}}\in[\kappa\setminus\eta]^{<\kappa}$ and $f_{\alpha^{\prime}}:L_{\alpha^{\prime}}\hookrightarrow\bigcup_{\beta\in K_{\alpha^{\prime}}}L_{\beta}\cap C$; the recursion works.
Let $f=\bigcup_{\alpha\in\kappa}f_{\alpha}$. By (ii) and (iii), $\alpha<\alpha^{\prime}$ implies $f_{\alpha}[L_{\alpha}]<f_{\alpha^{\prime}}[L_{\alpha^{\prime}}]$ and, hence, $f:L\hookrightarrow C$. Thus $C^{\prime}=f[L]\in{\mathbb{P}}(L)$, $C^{\prime}\subset C$ and, by Fact 4.3, $C\in{\mathbb{P}}(L)$.